Suppose General Surplus Corporation sells personal copying machines at the rate of 1000 units per year. The cost of placing one order is $180, and it costs $100 per year to carry a copier in inventory. Which of the below statements is false?

Group of answer choices

The number of orders per year is 17 (rounded to the nearest integer), and the time interval between orders is 22 days (rounded to the nearest integer) .

The average inventory is 30 copiers.

The annual carrying cost is $2,700.

The annual order cost is $3,000.
Answer #1

Option D The annual order cost is $3,000.

Economic Order Quantity=sqrt(2*Annual Demand*Ordering cost/Carrying Cost)=sqrt(2*1000*180/100)=60

Number of orers per year=1000/60=16.66667 or 17

Time interval=365/17=21.47059 or 22 days

Average inventory=60/2=30

Annual carrying cost=30*100=3000

Annual order cost=17*180=3060